# On the Robustness of Kernel Goodness-of-Fit Tests

### v4

@@ 1 Introduction
Goodness-of-fit (GOF) testing D’Agostino,, 1986; Lehmann and Romano,, 2022, Chapter 16 tackles the question of how well a given probabilistic model describes some observed data. More formally, given a model $P$ and observations drawn independently from some distribution $Q$, GOF testing compares the null hypothesis $H_{0}:Q=P$ against the alternative hypothesis $H_{1}:Q\neq P$. This process is fundamental to validating a model before it is used for predictions, decision-making, or providing probabilistic guarantees on important quantities of interest. Indeed, GOF testing is closely linked to the concept of statistical significance, which now permeates almost all areas of science and engineering.
@@
In this paper, we focus exclusively on kernel-based GOF tests (Chwialkowski et al.,, 2016; Liu et al.,, 2016). Unlike most classical alternatives such as likelihood-ratio tests (Hogg et al.,, 1977, Chapter 8.7) and Kolmogorov-Smirnov tests (Kolmogorov,, 1933), kernel GOF tests can be used for models whose density function is known only up to a multiplicative constant. This is a significant advantage since it enables their use for models which are beyond the reach of other tests, including many modern flexible density estimation models, energy-based models, large graphical models, and Bayesian posteriors. Kernel-based GOF tests achieve this through a tractable test statistic based on the score function (i.e., the gradient of the log-density), which is widely available since it can be evaluated without knowledge of the normalization constant of the density. This property has made kernel GOF tests popular and has led to numerous extensions specializing the approach to problems involving time-to-event data (Fernandez et al.,, 2020), discrete data (Yang et al.,, 2018), point processes (Yang et al.,, 2019), manifold data (Xu and Matsuda,, 2020), graphs (Xu and Reinert,, 2021), protein structures (Amin et al.,, 2023), and text documents of variable lengths (Baum et al.,, 2023).
@@
A significant drawback of GOF testing is the following conundrum: in almost all real-world applications, models are wrong, leading GOF tests to ultimately reject the null hypothesis as the sample size grows. However, there is often a difference between statistical significance from a GOF perspective, and practical relevance of the model for downstream tasks. For instance, data corruption may arise in signal processing due to sensor failures (Rizzoni and Min,, 1991; Sharma et al.,, 2010), in classification tasks due to mislabelled entities (Frénay and Verleysen,, 2013), and in radio systems due to impulsive noise (Blackard et al.,, 1993). In these cases, $H_{0}$ may be rejected even though $P$ is almost correct up to these mild perturbations, and as such could still be a useful model for predictions or providing probabilistic statements on quantities of interest.
@@
This conundrum has led to the development of robust tests (Huber,, 1965; LeCam,, 1973; Dabak and Johnson,, 1994; Fauß et al.,, 2021), which aim to control the Type-I error rate (the probability of falsely rejecting the null hypothesis) under a composite null hypothesis $H^{\mathrm{C}}_{0}:Q\in\mathcal{P}_{0}$, where $\mathcal{P}_{0}$ is a family of probability distributions, called the uncertainty set (Fauß et al.,, 2021). This uncertainty set is constructed to include $P$ and distributions similar to $P$ in some appropriate notion. By designing $\mathcal{P}_{0}$ to include potential contaminations, such as adversarial contaminations or outliers, the composite null $H^{\mathrm{C}}_{0}$ can now hold in realistic scenarios, thus avoiding the aforementioned conundrum and aligning statistical and practical significance.
@@
Unfortunately, robust kernel GOF tests are absent from the existing literature on kernel tests. The closest work is that of Sun and Zou, (2023); Schrab and Kim, (2024), who proposed robust tests using uncertainty sets defined by the Maximum Mean Discrepancy (MMD; Müller, 1997; Gretton et al., 2012). These tests require samples from both $Q$ and $P$. However, sampling from $P$ might be infeasible or require significant additional computational cost. The approximation of $P$ by finite samples could also reduce statistical efficiency.
We propose a novel class of robust kernel GOF tests based on the kernel Stein discrepancy (KSD). These tests are appropriate in the one-sample setting, where samples are available only from $Q$, and an unnormalized density is available for $P$. To guarantee robustness against mild perturbations, we make use of tilted kernels proposed in a related construction for robust parameter estimation in Barp et al., (2019); Matsubara et al., (2022). We then study qualitative and quantitative robustness, two notions of robustness which capture a test’s insensitivity to perturbations in a neighborhood of $P$, as formally introduced in Section˜2.3. Our contributions are summarized in Table˜1 and detailed below:
@@
    In Section˜3, we study the robustness properties of existing KSD tests. We show that KSD tests with stationary kernels are not necessarily robust under infinitesimal model deviation (which we call qualitative robustness; cf. Definition˜1), and never robust against fixed classes of misspecified models (called quantitative robustness; cf. Definition˜2). We then show that qualitative robustness can be guaranteed by using appropriately tilted kernels, but this is not sufficient to guarantee quantitative robustness.
@@ 2.1 The Kernel Stein Discrepancy
A key ingredient for kernel GOF tests is the kernel Stein discrepancy (KSD, Chwialkowski et al.,, 2016; Liu et al.,, 2016; Gorham and Mackey,, 2017; Oates et al.,, 2017). Throughout, we assume $P,Q\in\mathcal{P}(\mathbb{R}^{d})$, and $P$ admits a Lebesgue density function $p\in\mathcal{C}^{1}$ such that $p>0$ on $\mathbb{R}^{d}$. Let $k\in\mathcal{C}^{(1,1)}$ be a scalar-valued reproducing kernel with associated reproducing kernel Hilbert space (RKHS, Berlinet and Thomas-Agnan,, 2004) denoted by $\mathcal{H}_{k}$. The Langevin KSD gives a notion of discrepancy between $Q$ and $P$ and is defined as
@@
where $\mathcal{B}\coloneqq\{h=(h_{1},\ldots,h_{d}):\;h_{j}\in\mathcal{H}_{k}\textrm{ and }\sum_{j=1}^{d}\|h_{j}\|_{\mathcal{H}_{k}}^{2}\leq 1\}$ is the unit ball in the $d$-times Cartesian product $\mathcal{H}_{k}^{d}$ of the RKHS $\mathcal{H}_{k}$. The operator $\mathcal{A}_{p}$ is a linear operator called the Langevin Stein operator and it maps (suitably regular) vector-valued functions $f$ to scalar-valued ones via $(\mathcal{A}_{p}f)(\mathbf{x})\coloneqq\mathbf{s}_{p}(\mathbf{x})^{\top}f(\mathbf{x})+\nabla^{\top}f(\mathbf{x})$, where $\mathbf{s}_{p}(\mathbf{x})\coloneqq\nabla_{\mathbf{x}}\log p(\mathbf{x})$ is known as the (Hyvärinen) score function of $P$ (Hyvärinen,, 2005). Throughout this paper, we will shorten Langevin KSD to KSD for brevity, but note that other Stein operators can also be used to constructs KSDs; see Anastasiou et al., (2023) for a review. When $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}\|_{2}]<\infty$, the squared KSD admits the following double-integral form Barp et al.,, 2024, Corollary 1; Gorham and Mackey,, 2017
@@
whenever $D^{2}(Q,P)$ is well-defined. The function $u_{p}$ is also a reproducing kernel, known as the Stein (reproducing) kernel. The main advantage of the KSD is that it is computable even if $p$ is unnormalized. Indeed, $u_{p}$ depends on $p$ only through $\mathbf{s}_{p}$, which does not depend on the normalizing constant of $p$ (as the constant is cancelled due to the differentiation). The KSD can be straightforwardly estimated at $\mathcal{O}(n^{2})$ cost using a V-statistic estimator, which is non-negative and biased but consistent. It is formed by replacing $Q$ with its empirical version $Q_{n}$ based on independent observations $\mathbb{X}_{n}\coloneqq\{\mathbf{X}_{i}\}_{i=1}^{n}$ drawn from $Q$:
@@
where we have overloaded the notation by writing $D^{2}:(\mathbb{R}^{d})^{n}\to[0,\infty)$ as a test statistic. An alternative estimator is the U-statistic, which differs from (2) by summing only over disjoint pairs $i\neq j$. Although the U-statistic is unbiased, it can take negative values. Therefore, in this paper, we focus on V-statistics, leaving a discussion on how to extend our results to U-statistics to Appendix˜C.
When the kernel $k$ is characteristic, the KSD is $P$-separating, meaning that $D(Q,P)\geq 0$ with equality if and only if $Q=P$ for all $Q$ that finitely integrates $\|\mathbf{s}_{p}\|_{2}$; see, e.g., Barp et al., (2024, Theorem 3). Most characteristic kernels used in kernel GOF tests are sufficiently smooth stationary kernels of the form $k(\mathbf{x},\mathbf{x}^{\prime})=h(\mathbf{x}-\mathbf{x}^{\prime})$, where $h\in\mathcal{C}^{2}_{b}$ and $h(0)>0$. These include the squared-exponential kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\exp(-\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}/(2\lambda^{2}))$ and the inverse multi-quadric (IMQ) kernel $k(\mathbf{x},\mathbf{x}^{\prime})=(1+\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}/\lambda^{2})^{-b}$, where $b>0$, and $\lambda>0$ is a hyperparameter known as the bandwidth. The IMQ kernel is often preferred in practice since the resulting KSD has the desirable property of $P$-convergence control (Gorham and Mackey,, 2017; Barp et al.,, 2024). The bandwidth $\lambda$ also plays a crucial role in the performance of kernel-based tests (Reddi et al.,, 2015; Ramdas et al.,, 2015; Huang et al.,, 2023), and a median-heuristic (Fukumizu et al.,, 2009, Section 5) is often used to select its value in practice (Liu et al.,, 2016; Chwialkowski et al.,, 2016).
@@ 2.2 Kernel Goodness-of-Fit Testing
The KSD provides a natural quantity to construct a test statistic since, by the $P$-separation property of KSD, testing $H_{0}:Q=P$ is equivalent to testing whether $D(Q,P)=0$. Since the KSD is non-negative, A GOF test can therefore be constructed whereby $H_{0}$ is rejected for large values of the KSD estimate (Chwialkowski et al.,, 2016; Liu et al.,, 2016). KSD tests have been used for GOF testing with unnormalized models in a wide range of applications and data structures (Yang et al.,, 2018; Fernandez et al.,, 2020; Xu and Reinert,, 2021; Xu and Matsuda,, 2021; Amin et al.,, 2023). Extensions have also been developed to address its limitations, such as high computational cost (Jitkrittum et al.,, 2017; Huggins and Mackey,, 2018), difficulty in bandwidth selection (Schrab et al.,, 2022) and lack of test power against certain alternatives (Liu et al.,, 2023).
A significant practical challenge with KSD tests is determining an appropriate decision threshold. A valid threshold can be obtained by considering the quantiles of the null distribution of $D^{2}(\mathbb{X}_{n})$, but since this distribution is not usually tractable, a bootstrap procedure is often employed to estimate it. This is done by computing the empirical quantiles based on bootstrap samples of the form
@@
where $\mathbf{W}\coloneqq(W_{1},\ldots,W_{n})\sim\textrm{Multinomial}(n;1/n,\ldots,1/n)$. This procedure, called Efron’s bootstrap or weighted bootstrap (Arcones and Gine,, 1992; Janssen,, 1994), assigns a multinomial weight to each observation, which mimics recomputing the test statistic using data re-sampled from $\mathbb{X}_{n}$ with replacement. Other bootstrap methods such as wild bootstraps (Leucht and Neumann,, 2013; Shao,, 2010) are also viable, particularly when the samples are potentially correlated (Chwialkowski et al.,, 2016), but we focus on the weighted bootstrap since they perform similarly for independent samples, as shown in Section˜B.8. The test threshold is selected as the $(1-\alpha)$-quantile of the distribution of $D^{2}_{\mathbf{W}}(\mathbb{X}_{n})$ conditional on $\mathbb{X}_{n}$, i.e.,
@@
where $\mathbbm{1}\{\mathcal{A}\}$ denotes the indicator function for the event $\mathcal{A}$. The KSD test then rejects $H_{0}$ if $D^{2}(\mathbb{X}_{n})>q^{2}_{B,1-\alpha}(\mathbb{X}_{n})$. This test is asymptotically well-calibrated for $H_{0}:Q=P$ (Liu et al.,, 2016, Theorem 4.3). When KSD is $P$-separating, this test is also consistent, meaning that, whenever $Q\neq P$, the probability of rejection approaches one as the sample size $n\to\infty$ (Chwialkowski et al.,, 2016; Liu et al.,, 2016).

### v5

@@ Abstract
Keywords: robustness, hypothesis testing, kernel methods, Stein’s method
@@ 1 Introduction
Goodness-of-fit (GOF) testing D’Agostino, 1986; Lehmann and Romano, 2022, Chapter 16 tackles the question of how well a given probabilistic model describes some observed data. More formally, given a model $P$ and observations drawn independently from some distribution $Q$, GOF testing compares the null hypothesis $H_{0}:Q=P$ against the alternative hypothesis $H_{1}:Q\neq P$. This process is fundamental to validating a model before it is used for predictions, decision-making, or providing probabilistic guarantees on important quantities of interest. Indeed, GOF testing is closely linked to the concept of statistical significance, which now permeates almost all areas of science and engineering.
@@
In this paper, we focus exclusively on kernel-based GOF tests (Chwialkowski et al., 2016; Liu et al., 2016). Unlike most classical alternatives such as likelihood-ratio tests (Hogg et al., 1977, Chapter 8.7) and Kolmogorov-Smirnov tests (Kolmogorov, 1933), kernel GOF tests can be used for models whose density function is known only up to a multiplicative constant. This is a significant advantage since it enables their use for models which are beyond the reach of other tests, including many modern flexible density estimation models, energy-based models, large graphical models, and Bayesian posteriors. Kernel GOF tests achieve this through a tractable test statistic based on the score function (i.e., the gradient of the log-density), which is widely available since it can be evaluated without knowledge of the normalization constant of the density. This property has made kernel GOF tests popular and has led to numerous extensions specializing the approach to problems involving time-to-event data (Fernandez et al., 2020), discrete data (Yang et al., 2018), point processes (Yang et al., 2019), manifold data (Xu and Matsuda, 2020), graphs (Xu and Reinert, 2021), protein structures (Amin et al., 2023), and text documents of variable lengths (Baum et al., 2023).
@@
A significant drawback of GOF testing is the following conundrum: in almost all real-world applications, models are wrong, causing GOF tests to ultimately reject the null hypothesis as the sample size grows. However, there is often a difference between statistical significance from a GOF perspective, and the practical relevance of the model. For instance, data corruption may arise in signal processing due to sensor failures (Rizzoni and Min, 1991; Sharma et al., 2010), in classification tasks due to mislabelled entities (Frénay and Verleysen, 2013), and in radio systems due to impulsive noise (Blackard et al., 1993). In these cases, $H_{0}$ may be rejected even though $P$ is almost correct up to these mild perturbations, and thus it could still be useful for downstream tasks such as prediction or probabilistic inference.
@@
This conundrum has led to the development of robust tests (Huber, 1965; LeCam, 1973; Dabak and Johnson, 1994; Fauß et al., 2021), which aim to control the Type-I error rate (the probability of falsely rejecting the null hypothesis) under a composite null hypothesis $H^{\mathrm{C}}_{0}:Q\in\mathcal{P}_{0}$. The set $\mathcal{P}_{0}$ is a family of probability distributions, called the uncertainty set (Fauß et al., 2021), which is constructed to include $P$ and distributions similar to $P$ in some appropriate notion. By designing $\mathcal{P}_{0}$ to contain potential contaminations, such as adversarial contaminations or outliers, the composite null hypothesis $H^{\mathrm{C}}_{0}$ can now hold in realistic scenarios, thus avoiding the aforementioned conundrum and aligning statistical and practical significance.
@@
Unfortunately, robust kernel-based GOF tests have only received limited attention in the existing literature. The closest work is that of Sun and Zou (2023); Schrab and Kim (2024), who proposed robust tests using uncertainty sets defined by the Maximum Mean Discrepancy (MMD; Müller 1997; Gretton et al. 2012). These tests require samples from both $Q$ and $P$. However, generating samples from $P$ may be infeasible or computationally expensive, and approximating $P$ by finite samples can also reduce statistical efficiency.
We propose a novel class of robust kernel GOF tests based on the kernel Stein discrepancy (KSD). These tests are appropriate in the one-sample setting, where we assume samples are available only from $Q$, and an unnormalized density is available for $P$. To guarantee robustness against mild perturbations, we use tilted kernels proposed in Barp et al. (2019); Matsubara et al. (2022), in a related construction for robust parameter estimation. We then study qualitative and quantitative robustness, two notions of robustness which capture a test’s insensitivity to perturbations around $P$, as formally introduced in Section˜2.3. Our contributions are summarized in Table˜1 and detailed below:
@@
    In Section˜3, we study the robustness properties of existing KSD tests. We show that KSD tests with stationary kernels are not necessarily robust under infinitesimal model deviation (which we call qualitative robustness; see Definition˜1), and never robust against fixed classes of misspecified models (called quantitative robustness; see Definition˜2). We then show that qualitative robustness can be guaranteed by using appropriately tilted kernels, but this is not sufficient to guarantee quantitative robustness.
@@ 2.1 The Kernel Stein Discrepancy
A key ingredient for kernel GOF tests is the kernel Stein discrepancy (KSD, Chwialkowski et al., 2016; Liu et al., 2016; Gorham and Mackey, 2017; Oates et al., 2017). Throughout, we assume $P,Q\in\mathcal{P}(\mathbb{R}^{d})$, and $P$ admits a Lebesgue density function $p\in\mathcal{C}^{1}$ such that $p>0$ on $\mathbb{R}^{d}$. Let $k\in\mathcal{C}^{(1,1)}$ be a scalar-valued reproducing kernel with associated reproducing kernel Hilbert space (RKHS, Berlinet and Thomas-Agnan, 2004) denoted by $\mathcal{H}_{k}$. The Langevin KSD gives a notion of discrepancy between $Q$ and $P$ and is defined as
@@
where $\mathcal{B}\coloneqq\{h=(h_{1},\ldots,h_{d}):\;h_{j}\in\mathcal{H}_{k}\textrm{ and }\sum_{j=1}^{d}\|h_{j}\|_{\mathcal{H}_{k}}^{2}\leq 1\}$ is the unit ball in the $d$-times Cartesian product $\mathcal{H}_{k}^{d}$ of the RKHS $\mathcal{H}_{k}$. The operator $\mathcal{A}_{p}$ is a linear operator called the Langevin Stein operator and it maps (suitably regular) vector-valued functions $f$ to scalar-valued ones via $(\mathcal{A}_{p}f)(\mathbf{x})\coloneqq\mathbf{s}_{p}(\mathbf{x})^{\top}f(\mathbf{x})+\nabla^{\top}f(\mathbf{x})$, where $\mathbf{s}_{p}(\mathbf{x})\coloneqq\nabla_{\mathbf{x}}\log p(\mathbf{x})$ is known as the (Hyvärinen) score function of $P$ (Hyvärinen, 2005). Throughout this paper, we will shorten Langevin KSD to KSD for brevity, but note that other Stein operators can also be used to construct KSDs; see Anastasiou et al. (2023) for a review. When $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}\|_{2}]<\infty$, the squared KSD admits the following double-integral form Barp et al., 2024, Corollary 1; Gorham and Mackey, 2017
@@
whenever $D^{2}(Q,P)$ is well-defined. The function $u_{p}$ is also a reproducing kernel, known as the Stein (reproducing) kernel. The main advantage of the KSD is that it is computable even if $p$ is unnormalized. Indeed, $u_{p}$ depends on $p$ only through $\mathbf{s}_{p}$, which does not depend on the normalizing constant of $p$ (as the constant is cancelled due to the differentiation). The KSD can be straightforwardly estimated at $\mathcal{O}(n^{2})$ cost using a V-statistic estimator. It is formed by replacing $Q$ with its empirical version $Q_{n}$ based on independent observations $\mathbb{X}_{n}\coloneqq\{\mathbf{X}_{i}\}_{i=1}^{n}$ drawn from $Q$:
@@
where we have overloaded the notation by writing $D^{2}:(\mathbb{R}^{d})^{n}\to[0,\infty)$ as a test statistic. The V-statistic is non-negative, biased but consistent.
An alternative estimator is a U-statistic, which differs from (2) by summing only over disjoint index pairs $i\neq j$. Although a U-statistic is unbiased, it can take negative values. Therefore, we focus on V-statistics in this paper, and leave a discussion on how to extend our results to U-statistics to Appendix˜C.
When the kernel $k$ is characteristic, the KSD is $P$-separating, meaning that $D(Q,P)\geq 0$, with equality if and only if $Q=P$ for all $Q$ that finitely integrates $\|\mathbf{s}_{p}\|_{2}$; see, e.g., Barp et al. (2024, Theorem 3). Most characteristic kernels used in kernel GOF tests are sufficiently smooth stationary kernels of the form $k(\mathbf{x},\mathbf{x}^{\prime})=h(\mathbf{x}-\mathbf{x}^{\prime})$, where $h\in\mathcal{C}^{2}_{b}$ and $h(0)>0$. These include the squared-exponential kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\exp(-\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}/(2\lambda^{2}))$ and the inverse multi-quadric (IMQ) kernel $k(\mathbf{x},\mathbf{x}^{\prime})=(1+\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}/\lambda^{2})^{-b}$, where $b>0$, and $\lambda>0$ is a hyperparameter known as the bandwidth. The IMQ kernel is often preferred in practice since the resulting KSD has the desirable property of $P$-convergence control (Gorham and Mackey, 2017; Barp et al., 2024). The bandwidth $\lambda$ also plays a crucial role in the performance of kernel-based tests (Reddi et al., 2015; Ramdas et al., 2015; Huang et al., 2023), and a median-heuristic (Fukumizu et al., 2009, Section 5) is often used to select its value in practice (Liu et al., 2016; Chwialkowski et al., 2016).
@@ 2.2 Kernel Goodness-of-Fit Testing
KSD is a natural choice of test statistic for GOF testing, since, by the $P$-separation property of KSD, testing $H_{0}:Q=P$ is equivalent to testing whether $D(Q,P)=0$. As KSD is non-negative, A GOF test can therefore be constructed whereby $H_{0}$ is rejected for large values of the KSD estimate (Chwialkowski et al., 2016; Liu et al., 2016). KSD tests have been used for GOF testing with unnormalized models in a wide range of applications and data structures (Yang et al., 2018; Fernandez et al., 2020; Xu and Reinert, 2021; Xu and Matsuda, 2021; Amin et al., 2023). Extensions have also been developed to address its limitations, such as high computational cost (Jitkrittum et al., 2017; Huggins and Mackey, 2018), difficulty in bandwidth selection (Schrab et al., 2022) and lack of test power against certain alternatives (Liu et al., 2023).
A significant practical challenge with KSD tests is determining an appropriate decision threshold. A valid threshold can be obtained by considering the quantiles of the null distribution of $D^{2}(\mathbb{X}_{n})$, but since this distribution is not usually tractable, bootstrapping is often employed to estimate it. One approach is to compute the empirical quantiles of bootstrap samples of the form
@@
where $\mathbf{W}\coloneqq(W_{1},\ldots,W_{n})\sim\textrm{Multinomial}(n;1/n,\ldots,1/n)$. This procedure, called Efron’s bootstrap or weighted bootstrap (Arcones and Gine, 1992; Janssen, 1994), assigns a multinomial weight to each observation, which mimics recomputing KSD using data sampled with replacement from $\mathbb{X}_{n}$. Other bootstrap methods such as wild bootstraps (Leucht and Neumann, 2013; Shao, 2010) are also viable, particularly when the samples are potentially correlated (Chwialkowski et al., 2016), but we focus on the weighted bootstrap since we find that they perform similarly in our setting with independent samples; see Section˜B.8.
The test threshold is selected as the $(1-\alpha)$-quantile of the distribution of $D^{2}_{\mathbf{W}}(\mathbb{X}_{n})$ conditional on $\mathbb{X}_{n}$, i.e.,
@@
where $\mathbbm{1}\{\mathcal{A}\}$ denotes the indicator function for the event $\mathcal{A}$. The KSD test then rejects $H_{0}$ if $D^{2}(\mathbb{X}_{n})>q^{2}_{B,1-\alpha}(\mathbb{X}_{n})$. This test is asymptotically well-calibrated for $H_{0}:Q=P$ (Liu et al., 2016, Theorem 4.3). When KSD is $P$-separating, this test is also consistent, meaning that, whenever $Q\neq P$, the probability of rejection approaches one as the sample size $n\to\infty$ (Chwialkowski et al., 2016; Liu et al., 2016).